OpenStudy (anonymous):

Solve the equation for x. Do not include "x = " in your answer. If necessary, enter a non-integer result as a fraction using the slash ( / ) bar. the square root 4-3x divided by the square root 3x = 2

OpenStudy (anonymous):

guess this is an on line class, huh?

OpenStudy (anonymous):

\[\sqrt{4-3x}=2\sqrt{3x}\] is a start square both sides to get \[4-3x=12x\] and solve for \(x\)

OpenStudy (anonymous):

yayyy! thanks :)

OpenStudy (anonymous):

yw